What is the pressure in a 5.00 L tank with 3.75 moles of oxygen at 39.3 °C?
Answer #1

Given:
V = 5.0 L
n = 3.75 mol
T = 39.3 oC
= (39.3+273) K
= 312.3 K

use:
P * V = n*R*T
P * 5 L = 3.75 mol* 0.08206 atm.L/mol.K * 312.3 K
P = 19.2205 atm
Answer: 19.2 atm